elimination/ addtion method
x - 3y = 4
x + 5y = -4

Answer by solver15(27)

multipy -1 to the equation x-3y=4
your equation will be
-x+3y=-4
then elimate -x with x
-x+3y=-4
x+5y=-4
-x and x is now elimated, so the equation will be
8y=-8, the divide both sides by 8
8y/8=-8/8
then y=-1
plug in -1 to the previous equation
x-3(-1)=4
x+3=4, the subtract 3 form both sides of the equation
-3 -3
the x=1, so your solution is (1,-1)